UT Sep 05 2022 - Night #4 of Restart




Night Info

  • Sunset 18:40
  • Nautical Twilight Ends 19:35
  • Astro Twilight Ends 20:05
  • Astro Twilight Begins 04:33
  • Nautical Twilight Begins 05:03
  • Sunrise 05:57
  • Moonset 23:27
  • Moon illumination ~66%

Staff

  • OSA: D.G. Huerta
  • ISp: E.Solheid
  • Mtn Manager: B. Waggoner
  • ISA/SO: B.Rothberg, A.Becker, R.T.Gatto
  • Technical Expert: J.Hill

Night Plan

LUCI-1 is offline until MCE/ROE power supply is replaced (or other explanation/fix found; it was a loose power cord)

Starting configuration PEPSI-POL Bino

ALTA Predictied Seeing

20220904_see_evol_time_Height_20_20000.png

PEPSI Checkout Part Deux

  • DONE Checkout of AGW 3/4 for PEPSI-Pol
    • This is dependent on the results of work on ITs 8701 (AGW 4 issues) and 8702 (Pinhole adjustment for SX) - if not fixed will skip to AGW 1/2 checkout
    • Verify SH spots are well illuminated
    • Send Acquire Preset
    • Active Preset and test collimation
  • DONE PEPI-Pol Flux Calibration Star
  • Name V Inst Fiber BlueArm RedArm Start MST End Duration Phase Moon PID
    gamma Equ 4.70 POL 200P CD3 6x00:04:00 CD5 6x00:04:00 19:20:00 19:48:18 00:28:18 84 polstd

  • HR 7950 3.78 POL 200P CD3 2x00:01:00 CD5 2x00:01:00 19:50:18 19:53:44 00:03:26 71 fluxstd

    • Gamma Equ 21 10 20 +10 07 53 - visible until 02:45 local time
    • HR 7950 20 47 40 -09 29 44 - visible until 01:20 local time
  • Other On-Sky PEPSI/Pol?
  • Reconfigure for PEPSI-PFU - 0.25hr
  • Name V Inst Fiber BlueArm RedArm Start MST End Duration Phase Moon PID

  • HR 7950 3.78 PFU 100 CD3 00:03:00 CD5 00:03:00 19:55:44 20:00:06 00:04:22 70 fluxstd
    HR 7950 3.78 PFU 200 CD3 00:02:00 CD5 00:02:00 20:02:06 20:04:49 00:02:43 70 fluxstd
    HR 7950 3.78 PFU 300 CD3 00:00:40 CD5 00:04:00 20:06:49 20:11:08 00:04:19 70 fluxstd

    • HR 7950 20 47 40 -09 29 44 - visible until 01:20 local time

AGW 1 & 2 Checkout Continued

Telescope Wakeup

  • DONE SX M3 Alignment LFBG - Critical (3 hrs) - requires LUCI AGWs - both sides can be done in parallel - requires SH spot and hotspot check completed
  • DONE DX M3 Alignment RFBG - Critical (3 hrs) - requires LUCI AGWs - both sides can be done in parallel - requires SH spot and hotspot check completed
    • 1.2" or better seeing
    • Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
  • Field Aberration (2 hrs) - Critical (1hour) (SX more important than DX due to SX removal)
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute
    • must complete SH spot centering and hotspot check before executing
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LUCI Wakeup

  • As time permits and if LUCI-1 is avaible (IT 7583 - unable to communicate with MCE/ROE)
    • if time can proceed with LUCI-2 checkout
  • LUCI Scripts Location: Service_Observing/Restart_2022/LUCI
    • LUCI_2022_Restart_V1.74.xml - all scripts (except LS Rot ZP)
    • RotZP _v2.xml - for LS RotZP scripts
  • Reconfigure from PEPSI PFU to LUCIs (0.1 hr)
  • Rotator center
  • rotator zeropoint (requires MOS mask) - (0.5 hr)
    • 1.2" or better seeing needed
    • LUCI-1 is empty - use the LS method
    • LUCI-2 has old masks from Feb 2022
      • can use Stone O and Stone P - note that Stone P had issues and required removing 1-2 stars.
    • Masks:
      • ID 903033 - Stone O
      • ID 921934 - Stone A
      • iD 931027 - Stone D
      • Do the LS method Bino to compare L2 with mask and LS method
    • Do a MOS acquisition (imaging of mask and thru-mask image confimration) - verify the theta < 0.1 deg for each mask
    • To update results (if needed)
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Longslit Method (backup in case mask exchange(s) aren't completed in time
      • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
      • use pseudo-LMS files below for acquisition - works in similar manner to MOS acquisition - will need to manually load the LMS files in RTD
  • Collect transform data - (2hrs) - must do RotZP First
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
  • Imaging check (1hr)
    • ANY seeing OK
    • dither in-field
      • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am
    • dither out-of-field (run one)
      • M52 - JHK, dither 5' out of field - visible all night
      • M40 - Broad filters, few narrow filters - visible all night
  • Non-sidereal imaging check (1hr)
    • dither in-field
      • TBD
    • dither out-of-field
      • TBD

LBC Checkout Continued

  • in case we are super efficient or can't continue with LUCI/AGW1-2 checkouts
  • Superfocus sequence -PA=0,180, 90, 270 - high (1 hr)
    • Requires seeing 1.2" or better, some clouds (thicker clouds require brighter stars)
  • Standard photometric fields - medium - (2hr)
    • must be clear, any seeing
    • minimum two airmasses needed
  • Spectro-photometric standards - low - (1hr)
    • must be clear, any seeing
  • Test Science OBs
    • cycle through U-Bessell and other Bessell filters
    • any seeing
    • M31 21hr-twilight
    • M33 22:15hr-twilight
    • Non-sidereal (flexible conditions)
      • 2020 AN3 - visible all night
      • 2018 XG5 - end of night

Night Log

Closed Dome

  • UT 01:24 - bringing up LBCs for later use, starting with PEPSI-Pol Checkout
  • UT 01:48 - opening dome

PEPSI Pol Checkout Part Troi

  • UT 02:00 - pointing preset
  • UT 02:17 - TSS triggered while pointing star preset was in progress, moving off to another position to see if its disabled (had a spurious enable beforeopening)
  • UT 02:30 - moving to collimation star
  • UT 02:33 - alarm going off on summit? Flashing lights and alarms on polycom!
  • UT 02:33 - right primary mirror panic
  • UT 02:38 - false fire alarm
  • UT 02:41 - now trying to fix moirror panics
  • UT 02:55 - pointing preset (again)
  • winds too high moving to another target
  • UT 02:58 - moving to Alpha Boo
    • too bright, saturated AGW
  • UT 03:08 - moving to Kappa CrB
  • UT 03:10 - collimation issues on SX
    • AGW 4 collimation looks good, SH spots well illuminated, seeing ~0.8-0.9"
  • UT 03:48 - finished, presetting to gamma Equ
  • UT 04:17 - completed Gamma Equ
  • UT 04:18 - reconfiguring for bino-LUCIs
  • PEPSI-Pol is ready for science!

AGW1/AGW2 checks

  • UT 04:31 - those damn co-pointing limit errors again
  • UT 04:39 - sending acquire preset
  • already backuped up oacontrol.conf.20220904
  • UT 04:44 - active preset for collimation
  • Starting AGW1 hotspot
    • stopping for a pointing check
  • UT 04:52 - moving for a pointing check, DX is lost
  • UT 05:00 - still no joy on DX, moving to a pointing star
  • UT 05:04 - now got it on DX
  • UT 05:10 - collimation preset
  • Starting AGW1 hotspot rotator check
    • ptincrement 0 CA -10
    • [oac@oac etc]$ setxy -u 1 -x 0 -y -612.5
    • setxy: setxy() failed: AGW1: Positive limit switch on localhost
    • "this is bad"
    • moving to +500 to test
    • [oac@oac etc]$ setxy -u 1 -x 0 -y 500
    • Command successfully transmitted
    • BR fault - i put in a "-"
    • short 1sec exposure, setting slew to hold 360
    • taking exposure
    • setting up AGW2 now
    • back to math on AGW1
    • X=112, Y=79
    • x2 --> X=224, Y=158
    • correct fo DETSEC X+15, Y+39---> X=239, Y=197
    • Hotspot is (x,y): 249.00, 196.00
    • X +10, Y=-1
    • 1 pixel = 1/20 arcsec for Potsdam AGw; and 0.600 mm per arcsec ---> 10 * 0.05 = 0.5--> 0.3 mm
    • The ROT axis has approximately 68 counts/mm,
    • Decreasing ROT homeoffset moves the star image (center of rotation) toward larger X pixel values
    • we want -20.4
    • From oaccontrol.conf
    • # 24998 refined on-sky to 24978 20220905 BR
      #homeoffset=24998,&1Q40
      #homeoffset=10613,&1Q40
      homeoffset=24978,&1Q40
    • rdwrconfig -u 1
    • UT 05:30 - restart GCS left
    • the AGW2 image got saturated, not sure why
    • UT 05:31 - remeasure the hotspot on AGW1 then worry about AGW2
    • UT 05:34 - guidestar missed on left after new preset
    • X=119, Y=80
    • x2---> X=238, Y=169
    • detsec correct X+15, Y+39 --> X=253, Y=208
    • Hotspot is (x,y): 249.00, 196.00 --> leaving as is - circle was jumpy so Y is less confident but we moved the right way for X
  • AGW2 check
    • ptincrement 1 CA -10
    • taking 90sec exposure on right guidecam
    • there is no circles in the image!?
    • UT 05:48 - restoring CA values and re-sending active preset
    • trying again
    • taking 90 sec exposure after offset ptincreent 1 CA -10 and set -u 2 -x 0 -y 612.5
    • same thing again no circle just a saturated blob
    • measuring halo of blob
    • X=130, Y=112
    • x2 ---> X=260, Y=224
    • DETSEC correct X+15, Y+39 --> X=275, Y=263
    • Hotspot is (x,y): 270.00, 238.00 --> hmmm, i'm not confident in this measurement
    • undoing CA -10 (so ptincrement 1 CA 10)
    • resending preset on DX
    • stopping guiding, slew to hold 0
    • ptincrement 1 IE 5
    • setxy -u 2 -x 0 -y 612.5
    • throwaway exposure
    • taking 90sec exposure
    • X=128, Y=114
    • x2 --> X=256, Y=228
    • DETSEC correction X+15, Y+39 --> X=271 Y=267 Hotspot is (x,y): 270.00, 238.00
    • Y needs to move -29
    • OOFSET calculation --> 0.87 mm decresing moves to larger Y pixels, so we need to increase to reduce Y +0.87mm * 1000 = +870 microns
    • AGW2 AXIS_RAD oofset updated in oaccontrol.conf
    • rdwrconfig -u 2
    • UT 06:11 - restart GCS right
    • restored IE on DX
    • UT 06:12 - resending preset to verify
    • offsetting in IE +5
    • setxy -u 2 -x 0 -y 612.5
    • taking 90sec exposure, slew to hold 0
    • X=128, Y=100
    • x2 ---> X=256, Y=200
    • DETSEC correct X+15, Y+39 ----> X=271, Y=239 Hotspot is (x,y): 270.00, 238.00
    • Good enough!
    • AGW1/2 Rotator hotspot complete!
    • undid IE offset DX

M3 Alignment Pupil Wobble Data Collection

  • UT 06:20 - sending active preset to same star - collimate on-axis
  • UT 06:23 - starting pupil_collect
    • saved in data/20220905_lfbg and data/20220905_rfbg
    • % Compiled module: IRC_OFFSETPOINTING.
      % Compiled module: COLLIMATE_ONCE.
      % Compiled module: GCS_TAKE_IMAGE.
      Acquiring RIGHT WFSC image (exp=15.0000 s)...
      % Attempt to subscript RESULT with <INT ( 2)> is out of range.
      % Execution halted at: GCS_TAKE_IMAGE 170 /home/lbto/idl/wfsc/gcs/gcs_take_image.pro
      % COLLIMATE_ONCE 53 /home/lbto/idl/wfsc/collimate/collimate_once.pro
      % PUPIL_COLLECT 70 /home/lbto/idl/wfsc/pupil/pupil_collect.pro
      % $MAIN$
    • BR error didn't press STOP WFS'ing button in GCSGUI
    • Working OK now, DIMM is 1.2" guiders were 0.8" as we started
  • UT 06:36 done
  • /lbt/lbto/supportscripts/MetrologySupport/m3grace.py -->need to add in bashrc later
  • Can't stat file ../lfbg_pupil_wobble.par
    Unable to read parameter file ../lfbg_pupil_wobble.par
  • above needs to be fixed, J.Hill is working on this now
  • use /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p [tip] -t [tilt]
  • PSFGUI left --> tertiay
    • Initial Values:
    • Global offsets are 0
    • TIP = -90, Tilt = -319.9, RZ (selector angle) = 28.0
    • moving +50 in tip /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p 50 -t 0
    • TIP = -40, Tilt = -320, RZ=28.0
  • UT 06:46 - running pupil_collect again on LFBG
  • PW_SX_take2.png
    • Undo the +50 in tip, and put in -50
    • /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p -50 -t -100
    • Current values are tip=-90, tilt-420, RZ=28
  • UT 07:00 - resend active preset
  • UT 07:03 - running pupil_collect again on LFBG
    • seeing is 1.5"-1.7" on guider, 1.7" on DIMM
  • Result
  • PW_SX_take3.png
    • change: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p -50 -t 200
    • and /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p 50 -t 0 since i up arrowed and left tip in there
    • Current values: tip -90, tilt -219, RZ=28
  • UT 07:15 - running pupil_collect on LFBG
  • RIGHT SIDE RESULTS (ignore the plot saying LFBG)
  • PW_DX_take1.png
    • Current DX Values: tip -160, tilt +120, RZ -25.0
    • change to /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p 50 -t 0
  • UT 07:21 - running pupil_collect DX
  • SX Round 4 results
  • PW_SX_take4.png
  • adding +100 in tilt /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p 0 -t 100
    • current values: tip -90, tilt -120, RZ=28
  • UT 07:29 - starting LFBG pupil_collect
  • RFBG Results
  • PW_DX_take2.png
    • adding +50 tip /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p 50 -t 0
    • Current Values: tip -60, tilt 120, RZ -25.7
  • UT 07:37 - starting DX pupil_collect
  • SX results, plotting the last two iterations only
  • PW_SX_take5.png
    • SX adding +50 tip: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p 50 -t 0
    • Current Values: Tip -40, Tilt-120, RZ=28
  • UT 07:51 - staring SX pupil_collect
  • DX Results Pass 3
  • PW_DX_take3.png
    • putting in -50 tip and -50 tilt: Tip -90, /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p -50 -t -50
    • Current Values: Tip -90, Tilt 109, RZ=-25.8
    • Got an error on moving M3
    • Making a DX M3 collimation update ....
      Error: right tertiary: Following error

      ptincrement 1 IE 0.670103683654308 ;ptincrement 1 CA 4.962672822882856
    • Last values were: tip -60, tilt 120, RZ -25.7 ---> only moved -30 in tip and +12 in tilt (this was wrong direction)
    • Lost DX guidestar
    • DX guidestar came back, tip and tilt values on PSFGUI right Tertiary just changed
    • J.Hill moved the tertiary to the correct values of Tip -110, Tilt 70, RZ=-25.8 BUT the guidestar is gone
  • UT 10 08:02 - B.Rothberg sending preset back to same target via IDL! BR taking the helm in a pinch, Mr. Sulu would approve....
    • iif_preset,PT10m094,mode=active,gs=0,rs=0,pa=0,/right
  • SX Complete - made it worse (plotting the last three iterations only)
  • PW_SX_take6.png
    • add +50 tilt: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p 0 -t 100
    • Current Values: Tip -40, Tilt -20, RZ = 28
  • UT 08:08 - starting pupil_collect on SX
  • +100 tilt was not correct
  • DX Results (only plotting last 3 takes)
  • PW_DX_take4.png
    • changing tilt -50: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p 0 -t -50
    • Current Values; Tip -109, Tilt 20, RZ -25.8
  • UT 08:20 - pupil_collect on DX
  • SX Results (plotting last 3 measurements)
  • PW_SX_take7.png
    • Adjust -50 tip and -50 tilt. /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p -100 -t -50
    • Current Values: Tilt -140, Tip -70, RZ 28.0
  • UT 08:25 - SX pupil_collect started
  • Guiders show seeing ~1.5", DIMM shows 1.5"-1.8"
  • DX Results (last three iterations plotted)
  • PW_DX_take5.png
    • adding +150 in tilt: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p 0 -t 150
    • Current values: Tip -110, Tilt 170, RZ -25.8
  • UT 08:37 - starting DX pupil collect
  • SX Results
  • PW_SX_take8.png
    • sending tip -50, tilt 0
    • Current Values: Tip -190, Tilt -70, RZ 28.0
  • UT 08:40 - starting SX pupil_collect
  • UT 08:42 - seeing is really getting bad - over 2" on guiders and DIMM
  • DX Results (all so far)
  • PW_DX_take6.png
    • sending tip -50, tilt 0: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p -50 -t 0
    • current values: Tip -160, Tilt 170, RZ -25.8
    • tweak: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s DX -p 0 -t -100
    • Current Values; Tip -60, Tilt 70, RZ -25.8
  • UT 08:54 - starting DX pupil_collect
  • SX Results (last 3 iterations)
  • PW_SX_take9.png
    • tweak: /lbt/lbto/supportscripts/MetrologySupport/m3tt.py -s SX -p 0 -t -50
    • Current Values: Tip -190, Tilt -120, RZ 28
  • UT 08:58 - starting SX pupil_collect
  • DX Results (last 4 plotted)
  • PW_DX_take7.png
    • This looks to be the best: DX Values are Tip = -160, Tilt = 70, RZ = -25.8 - Consider this DONE ADOPTED
  • SX Results (last 3 plotted)
  • PW_SX_take10.png
    • adding back 50 in tilt (this would correspond to the red data in the above plot)
    • SX Current Values: Tip = -190, Tilt = -70, RZ=28 ADOPTED
  • UT 09:11 - finished M3 Alignment - J.Hill will edit the M3 alignment config files in the afternoon
  • UT 09:12 - Reconfigure for Bino LBCs

LBC Checkout Continued

  • Check U-Bessell, collect some data
  • Red Pressure dropped out, will need to do lbckill / lbcstart - no alert, emailed instruments email list
2020 AN3 Observation
  • Going to be tight, might only get a few exposures
  • UT 09:39 - two co-pointing errors, PCS restarrted, still no joy
  • Mon Sep 5 09:39:29.409 2022 both Preset alert for LBC mode TRACK: PCS failed StartTrackingMount on side left because ERROR X Polynomial position is less than minimum limit. Pos. = -1.584582e+00 rad. Sl.
    Mon Sep 5 09:39:29.461 2022 both Preset alert for LBC mode TRACK: PCS StartTracking failed on side both with DEFAULT.
  • Even sending an IRAF preset to the coordinates doesn't work
  • Mon Sep 5 09:45:08.287 2022 left Preset alert for LBC mode TRACK: PCS failed StartTrackingMount on side left because ERROR X Polynomial position is less than minimum limit. Pos. = -1.571174e+00 rad. Sl.
    Mon Sep 5 09:45:08.367 2022 right Preset alert for LBC mode TRACK: PCS failed StartTrackingMount on side right because ERROR X Polynomial position is less than minimum limit. Pos. = -1.571173e+00 rad. Sl.
  • UT 09:47 - sending Copointing OB for collimation
  • UT 09:55 - Collimated sort of , DIMM reports 2.2"-3", moving to co-pointing field
  • UT 09:58 - sendnig NS script to target - might get one exposure
  • UT 10:00 - FW2 failed on BLUE
    • 35 2022/09/05 09:58:32.105493 FILTERS WHEEL#1 filter is more than 10 times the maximum allowed distance from its nominal position [src/filters/filters.c:2356]
      34 2022/09/05 09:58:27.802770 FILTERS WHEEL#1 timeout error on motion, took 19.524769 secs [src/filters/filters.c:2244]
      4 2022/09/05 09:58:53.723812 OB execution error: one or more systems failed [src/lbc/actions.c:1162]
      53 2022/09/05 09:58:36.910130 FILTERS WHEEL#2 filter is more than 10 times the maximum allowed distance from its nominal position [src/filters/filters.c:2356]
      52 2022/09/05 09:58:32.580923 FILTERS WHEEL#2 timeout error on motion, took 24.302956 secs [src/filters/filters.c:2244]
      51 2022/09/05 09:58:32.105818 >>> preset/expose/save interrupted at preset time, retcode=7
    • will go to zenith and power off and on LBC-Blue
  • didn't save so no data taken
  • UT 10:05 - at zenith - takig a 1sec test exposure U-Bessell LBCB and R-Bessell LBCR - did this by making a new OB in the GUI
    • 022/09/05 10:06:22.534231 OB execution error: one or more systems failed [src/lbc/actions.c:1162]
      46 2022/09/05 10:05:52.111701 >>> preset/expose/save interrupted at preset time, retcode=7
      45 2022/09/05 10:05:52.111615 filters system encountered a problem 13 [src/channel/channel.c:1157]
      44 2022/09/05 10:05:37.958367 FILTERS WHEEL#1 hardware failure [src/filters/filters.c:1485]
      43 2022/09/05 10:05:36.596701 FILTERS WHEEL#2 timeout error on motion, took 15.438761 secs [src/filters/filters.c:2244]
      42 2022/09/05 10:05:36.154582 FILTERS WHEEL#1 PositionWheel motion stop return error(3) or bad answer(p ) expecting v [src/filters/filters.c:2326], retry
      41 2022/09/05 10:05:35.308955 FILTERS WHEEL#1 timeout error on motion, took 14.150987 secs [src/filters/filters.c:2244]
  • UT 10:08 - powering off LBCs - similar to IT 5065 - powering on LBCs, will move to a new target
NGC 281 Test Filters
  • UT 10:15 - sending co-pointing OB for NGC 281 to collimate
  • UT 10:17 - running DOFPIA
  • UT 10:20 - updated IT 5065, seeing 1.5" on DIMM
  • UT 10:23 - copointing
  • UT 10:26 - starting science-like OB
  • UT 10:30 - a lot of noise in lbcr 102906, its a 1sec V-band image
  • UT 10:49 -seeing is all over the place, one minute its 1.2" on the DIMM and 1" on LBC images, next its 1.8" on the DIMM and 1.5" on LBC images
  • UT 10:59 - stopping after 5/7 dithers - will move to next target
    • lbcb.20220905.102621.fits-lbcb.20220905.110014.fits
    • lbcr.20220905.102512.fits - lbcr.20220905.105801.fits
NS 2018 XG5 Target
  • UT 11:03 - sending co-pointing script for collimation
  • UT 11:12 - copointing OB
  • UT11:15 - sending NS OB
  • UT 11:21 - original script had 5 exp per dither, took first exposure then made new OB with 1 NEXP per dither position
  • UT 11:27 - got 2exposures each filter before TSS got triggered - moving off
LBC Photometric Field (03hr Field)
  • UT 11:32 - Copointing OB sent for collimation
  • UT 11:38 - copointing
  • UT 11:39 - starting PhotField OB - Airmass 1.019
  • UT 12:03 - OB finished

LBC Skyflat

  • LBCB-V, LBCR-V PA=0
  • LBCB-V, LBCR-V PA=180
  • LBCB-U, LBCR-Y PA=0
  • LBCB-U, LBCR-Y PA=180 - got two exposures

Closed Dome

  • Taking 25 biases for LBCs

-- BarryRothberg - 04 Sep 2022
I Attachment Action Size Date Who Comment
20220904_see_evol_time_Height_20_20000.pngpng 20220904_see_evol_time_Height_20_20000.png manage 120 K 04 Sep 2022 - 21:21 BarryRothberg  
PW_DX_take1.pngpng PW_DX_take1.png manage 24 K 05 Sep 2022 - 07:18 BarryRothberg  
PW_DX_take2.pngpng PW_DX_take2.png manage 26 K 05 Sep 2022 - 07:34 BarryRothberg  
PW_DX_take3.pngpng PW_DX_take3.png manage 29 K 05 Sep 2022 - 07:52 BarryRothberg  
PW_DX_take4.pngpng PW_DX_take4.png manage 31 K 05 Sep 2022 - 08:17 BarryRothberg  
PW_DX_take5.pngpng PW_DX_take5.png manage 29 K 05 Sep 2022 - 08:31 BarryRothberg  
PW_DX_take6.pngpng PW_DX_take6.png manage 36 K 05 Sep 2022 - 08:52 BarryRothberg  
PW_DX_take7.pngpng PW_DX_take7.png manage 31 K 05 Sep 2022 - 09:06 BarryRothberg  
PW_SX_take10.pngpng PW_SX_take10.png manage 30 K 05 Sep 2022 - 09:10 BarryRothberg  
PW_SX_take2.pngpng PW_SX_take2.png manage 28 K 05 Sep 2022 - 06:57 BarryRothberg  
PW_SX_take3.pngpng PW_SX_take3.png manage 33 K 05 Sep 2022 - 07:13 BarryRothberg  
PW_SX_take4.pngpng PW_SX_take4.png manage 36 K 05 Sep 2022 - 07:26 BarryRothberg  
PW_SX_take5.pngpng PW_SX_take5.png manage 28 K 05 Sep 2022 - 07:41 BarryRothberg  
PW_SX_take6.pngpng PW_SX_take6.png manage 29 K 05 Sep 2022 - 08:06 BarryRothberg  
PW_SX_take7.pngpng PW_SX_take7.png manage 29 K 05 Sep 2022 - 08:21 BarryRothberg  
PW_SX_take8.pngpng PW_SX_take8.png manage 29 K 05 Sep 2022 - 08:38 BarryRothberg  
PW_SX_take9.pngpng PW_SX_take9.png manage 30 K 05 Sep 2022 - 08:56 BarryRothberg  
Topic revision: r6 - 06 Sep 2022, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback